# Self-Supervised Pre-training with Symmetric Superimposition Modeling for Scene Text Recognition

### v1

@@ Abstract
In text recognition, self-supervised pre-training emerges as a good solution to reduce dependence on expansive annotated real data. Previous studies primarily focus on local visual representation by leveraging mask image modeling or sequence contrastive learning. However, they omit modeling the linguistic information in text images, which is crucial for recognizing text. To simultaneously capture local character features and linguistic information in visual space, we propose Symmetric Superimposition Modeling (SSM). The objective of SSM is to reconstruct the direction-specific pixel and feature signals from the symmetrically superimposed input. Specifically, we add the original image with its inverted views to create the symmetrically superimposed inputs. At the pixel level, we reconstruct the original and inverted images to capture character shapes and texture-level linguistic context. At the feature level, we reconstruct the feature of the same original image and inverted image with different augmentations to model the semantic-level linguistic context and the local character discrimination. In our design, we disrupt the character shape and linguistic rules. Consequently, the dual-level reconstruction facilitates understanding character shapes and linguistic information from the perspective of visual texture and feature semantics. Experiments on various text recognition benchmarks demonstrate the effectiveness and generality of SSM, with 4.1% average performance gains and 86.6% new state-of-the-art average word accuracy on Union14M benchmarks.
@@ 2.2 Self-Supervised Text Recognition
Self-supervised learning technologies have been widely used for various works such as ML-LMCL Cheng et al. (2023b), TKDF Cheng et al. (2023c) and MESM Liu et al. (2024). The most representative paradigms for computer vision are Contrastive Learning Chen et al. (2021) and Masked Image Modeling (MIM) He et al. (2022); Bao et al. (2022); Chen et al. (2023). Recently, these self-supervised methods have been adopted for text recognition. SeqCLR Aberdam et al. (2021) is the pioneering work that proposes to model Contrast Learning on high-level sequence features for the first time. PerSec Liu et al. (2022) further performs Contrast Learning(CL) on both stroke-level features in shallow layers and Semantic-level features in deep layers. Later, DiG Yang et al. (2022) proposes to learn discriminative and generative features by integrating the CL and MIM. To explicitly focus on character structures and ensure sequence consistency, Guan proposes a character-level self-distillation framework Guan et al. (2023) based on unsupervised text segmentation maps. Recently, Union14M-UJiang et al. (2023), a 10M scale unlabeled real data is been presented together with MAERec-S, which uses MAE for self-supervised text recognition. In addition, MaskOCR Lyu et al. (2022) and DARLING Zhang et al. (2024) use synthetic data for text recognition pre-training.
Figure 2: The pre-training framework of SSM. The blue arrow and green arrow stand for the workflow of the student branch and teacher branch respectively. Origin View: original image, HFlip View: horizontally flipped image, VFlip View: vertically flipped image, Rotate View: 180-degree rotated image. $T_{P}$ and $T_{n}$ correspond to the original and the reversed text direction, respectively.
@@ 3 Methodology
The architecture of our proposed SSM is shown in Fig. 2, which comprises a Symmetric Superimposed Input Constructor, a student branch (blue arrow), and a teacher branch (orange arrow). We use the superimposed image of the original image and the horizontally flipped image as the input case to illustrate the whole workflow of our SSM.
@@ 3.1 Symmetrically Superimposed Input
Data Augmentation The superimposed image $X_{S}$ undergoes weak augmentations (e.g., gaussian blur, and grayscale conversion) to create a regular view $X^{reg}_{S}$ as the final input of student branch. The $X$ and $X_{R}$ are also transformed to $X^{reg}$ and $X^{reg}_{R}$ to supervise the image reconstruction of the student branch. For the teacher branch, we utilize the combination of both weak augmentations and geometry-based augmentations (e.g., affine transformation and perspective warping) to generate a pair of irregular views: $X^{irr}_{S}$, $X^{irr}$ and $X^{irr}_{R}$.
@@ 3.2 Symmetric Superimposition Modeling
2) Feature-level Representation Reconstruction To enhance the feature discriminability of character semantics and to model the spatial context between character visual semantics, we apply feature reconstruction in high-dimensional space. Specifically, we add the projector $\mathcal{H(\cdot)}$ followed by the Regressor $\mathcal{R(\cdot)}$ in the online branch. $\mathcal{H(\cdot)}$ aggregates the decoupled features ($F_{p}$ and $F_{n}$) into window-level features by adaptive mean-pooling and then maps them into high-level feature space, forming the corresponding prediction feature queries ($Q_{p}$ and $Q_{n}$). To construct the target representation, we introduce a target branch with the same structure as the student branch, except for the pixel decoder. $X^{irr}_{S}$ is fed into the target branch to obtain the window-level target feature keys ($K_{p}$ and $K_{n}$) on the fly. Next, we jointly employ the discriminative consistency loss $L_{dis}$ and dense reconstruction loss $L_{den}$ to supervise the reconstruction from $Q_{p}$ to $K_{p}$ and $Q_{n}$ to $K_{n}$ across the regular and irregular views. The loss $L_{dis}$ is to enhance the character classification representation at the semantic level, which can be formulated as:
@@
Since SiameseMIM Tao et al. (2023) demonstrates that predicting dense representations helps improve sensitivity to global structures, we directly minimize the distance between the predicted and the target features by $L_{mse}$ loss, aiming to better achieve semantic-level context modeling. Assuming that predictions are $Q=\{q^{i}\in\mathbb{R}^{d}|i=1,...,\mathcal{N}\}$ and targets is $K=\{k^{i}\in\mathbb{R}^{d}|i=1,...,\mathcal{N}\}$, where $\mathcal{N}$ is the total number of feature instances in one batch. The dense reconstruction loss $L_{den}$ ($L_{mse}$) between the original Q-K pairs ($Q_{p}$ and $K_{p}$) as well as inverted Q-K pairs ($Q_{n}$ and $K_{n}$) can be formulated as:
@@ 4.2 Implementation Details
Evaluation on Union14M benchmarks. Considering the saturation of the common benchmarks, we further compare SSM with other self-supervised methods on more challenging benchmarks in Tab. 2. Although we established a strong baseline model Scratch-ViT-Small, and SSM-ViT-Small still achieved an average performance gain of 3.5% and 2.6% on Union14M and other challenging data, respectively. Additionally, SSM surpassed MAE and DiG by 1.8% and 1.3% of average word accuracy on the Union14M benchmarks. The performance gains on all of Union14M benchmarks’ subsets demonstrate that SSM effectively learns intrinsic feature representations for different fonts, text orientations, and complex scenes. Surprisingly, SSM shows superior performance for contextless text images, which may be attributed to that embedding linguistic learning in visual space can avoid the text distribution dependence of explicit text correction.
@@ 4.4 Ablations and analysis
Effectiveness of architecture. The third row of Tab. 4 indicates that there was a 3.0% average performance improvement on Union14M benchmarks with the pixel-level reconstruction, which need the combined effect of prompt generator $\mathcal{G(\cdot)}$ transformer regressor $\mathcal{R(\cdot)}$ and pixel decoder $\mathcal{D(\cdot)}$. However, removing the regressor $\mathcal{R(\cdot)}$ results in a slight performance downgrade. The result demonstrates that the transformer regressor plays an indispensable role in SSM by guiding and disentangling the features of the original and inverted images. Based on the pixel reconstruction, the Joint use of the EMA mechanism and project module $\mathcal{H(\cdot)}$ further brings accuracy gains of 1.3% as it promotes learning discriminative character semantics and modeling spatial context at the semantic level. Here, $\mathcal{H(\cdot)}$ is mainly used to prevent excessive background noise in negative sample sampling and to map the decoupling feature to high-dimensional space. Lastly, the addition of Irregular view augmentations improves the average test accuracy up to 84.3% due to diverse perspectives.
Ablations on the types of superimposed input We compared our superposition strategies with add noise &\&& blur and randomly add another text image. As shown in Tab. 5(a), the performance gains of add another text image are close to that of add noise &\&& blur, both of which are lower than existing superposition strategies. The results can be attributed to that the model can easily reconstruct the two images with inconsistent text content and different background styles from low-level pixel differences, limiting the exploration of intrinsic character semantics and linguistic information. We also found that individually superposing a 180-degree rotation view is more effective than individually superposing a specific flip view. The result suggests that the rotation influences both character orientation and sequence reading order, helping capture learning richer character semantics and spatial context relationships during reconstruction. Finally, the best performance (84.3%) is achieved by combining HS, VS, and RS strategies, which improve the diversity of the superimposed input.
@@
Ablations on the feature-level reconstruction loss The first and second rows of Tab. 5(b) shows that only using distance measure functions such as MSE $L_{mse}$ and Cosine $L_{cos}$ results in lower word accuracy (82.5% vs 94.3%). We attribute this to that only minimizing the feature distance between predictions and targets may lead to feature collapse. Besides, Tab. 5(b) shows that our discriminative consistency loss $L_{dis}$ outperforms the dino loss $L_{dino}$ by 0.3%. The combined use of $L_{dis}$ and $L_{mse}$ ultimately yielded the best results. $L_{dis}$ can mitigate the feature collapse of $L_{mse}$ by reducing the similarity between negative samples. Besides, $L_{dis}$ focuses on enhancing discriminative features for character semantics while $L_{mse}$ is more sensitive to spatial structures, which helps model linguistic information in spatial context at the semantic level.

### v2

@@ Abstract
In text recognition, self-supervised pre-training emerges as a good solution to reduce dependence on expansive annotated real data. Previous studies primarily focus on local visual representation by leveraging mask image modeling or sequence contrastive learning. However, they omit modeling the linguistic information in text images, which is crucial for recognizing text. To simultaneously capture local character features and linguistic information in visual space, we propose Symmetric Superimposition Modeling (SSM). The objective of SSM is to reconstruct the direction-specific pixel and feature signals from the symmetrically superimposed input. Specifically, we add the original image with its inverted views to create the symmetrically superimposed inputs. At the pixel level, we reconstruct the original and inverted images to capture character shapes and texture-level linguistic context. At the feature level, we reconstruct the feature of the same original image and inverted image with different augmentations to model the semantic-level linguistic context and the local character discrimination. In our design, we disrupt the character shape and linguistic rules. Consequently, the dual-level reconstruction facilitates understanding character shapes and linguistic information from the perspective of visual texture and feature semantics. Experiments on various text recognition benchmarks demonstrate the effectiveness and generality of SSM, with 4.1% average performance gains and 86.6% new state-of-the-art average word accuracy on Union14M benchmarks. The code is available at https://github.com/FaltingsA/SSM.
@@ 2.2 Self-Supervised Text Recognition
Self-supervised learning technologies have been widely used for various works such as ML-LMCL Cheng et al. (2023b), TKDF Cheng et al. (2023c) and MESM Liu et al. (2024). The most representative paradigms for computer vision are Contrastive Learning Chen et al. (2021) and Masked Image Modeling (MIM) He et al. (2022); Bao et al. (2022); Chen et al. (2023). Recently, these self-supervised methods have been adopted for text recognition. SeqCLR Aberdam et al. (2021) is the pioneering work that proposes to model Contrast Learning on high-level sequence features for the first time. PerSec Liu et al. (2022) further performs Contrast Learning(CL) on both stroke-level features in shallow layers and Semantic-level features in deep layers. Later, DiG Yang et al. (2022) proposes to learn discriminative and generative features by integrating the CL and MIM. To explicitly focus on character structures and ensure sequence consistency, Guan proposes a character-level self-distillation framework Guan et al. (2023) based on unsupervised text segmentation maps. Recently, Union14M-UJiang et al. (2023), a 10M scale unlabeled real data has been presented together with MAERec-S, which uses MAE for self-supervised text recognition. In addition, MaskOCR Lyu et al. (2022) and DARLING Zhang et al. (2024) use synthetic data for text recognition pre-training.
Figure 2: The pre-training framework of SSM. The blue arrow and green arrow stand for the workflow of the online branch and target branch respectively. Origin View: original image, HFlip View: horizontally flipped image, VFlip View: vertically flipped image, Rotate View: 180-degree rotated image. $T_{P}$ and $T_{n}$ correspond to the original and the reversed text direction, respectively.
@@ 3 Methodology
The architecture of our proposed SSM is shown in Fig. 2, which comprises a Symmetric Superimposed Input Constructor, an online branch (blue arrow), and a target branch (orange arrow). We use the superimposed image of the original image and the horizontally flipped image as the input case to illustrate the whole workflow of our SSM.
@@ 3.1 Symmetrically Superimposed Input
Data Augmentation The superimposed image $X_{S}$ undergoes weak augmentations (e.g., gaussian blur, and grayscale conversion) to create a regular view $X^{reg}_{S}$ as the final input of online branch. The $X$ and $X_{R}$ are also transformed to $X^{reg}$ and $X^{reg}_{R}$ to supervise the image reconstruction of the online branch. For the target branch, we utilize the combination of both weak augmentations and geometry-based augmentations (e.g., affine transformation and perspective warping) to generate a pair of irregular views: $X^{irr}_{S}$, $X^{irr}$ and $X^{irr}_{R}$.
@@ 3.2 Symmetric Superimposition Modeling
2) Feature-level Representation Reconstruction To enhance the feature discriminability of character semantics and to model the spatial context between character visual semantics, we apply feature reconstruction in high-dimensional space. Specifically, we add the projector $\mathcal{H(\cdot)}$ followed by the Regressor $\mathcal{R(\cdot)}$ in the online branch. $\mathcal{H(\cdot)}$ aggregates the decoupled features ($F_{p}$ and $F_{n}$) into window-level features by adaptive mean-pooling and then maps them into high-level feature space, forming the corresponding prediction feature queries ($Q_{p}$ and $Q_{n}$). To construct the target representation, we introduce a target branch with the same structure as the online branch, except for the pixel decoder. $X^{irr}_{S}$ is fed into the target branch to obtain the window-level target feature keys ($K_{p}$ and $K_{n}$) on the fly. Next, we jointly employ the discriminative consistency loss $L_{dis}$ and dense reconstruction loss $L_{den}$ to supervise the reconstruction from $Q_{p}$ to $K_{p}$ and $Q_{n}$ to $K_{n}$ across the regular and irregular views. The loss $L_{dis}$ is to enhance the character classification representation at the semantic level, which can be formulated as:
@@
Since SiameseMIM Tao et al. (2023) demonstrates that predicting dense representations helps improve sensitivity to global structures, we directly minimize the distance between the predicted and the target features by $L_{mse}$ loss, aiming to achieve semantic-level context modeling. Assuming that predictions are $Q=\{q^{i}\in\mathbb{R}^{d}|i=1,...,\mathcal{N}\}$ and targets is $K=\{k^{i}\in\mathbb{R}^{d}|i=1,...,\mathcal{N}\}$, where $\mathcal{N}$ is the total number of feature instances in one batch. The dense reconstruction loss $L_{den}$ ($L_{mse}$) between the original Q-K pairs ($Q_{p}$ and $K_{p}$) as well as inverted Q-K pairs ($Q_{n}$ and $K_{n}$) can be formulated as:
@@ 4.2 Implementation Details
Evaluation on Union14M benchmarks. Considering the saturation of the common benchmarks, we further compare SSM with other self-supervised methods on more challenging benchmarks in Tab. 2. Although we established a strong baseline model Scratch-ViT-Small, and SSM-ViT-Small still achieved an average performance gain of 3.5% and 2.6% on Union14M and other challenging data, respectively. Additionally, SSM surpassed MAE and DiG by 1.8% and 1.3% of average word accuracy on the Union14M benchmarks. The performance gains on all of Union14M benchmarks’ subsets demonstrate that SSM effectively learns intrinsic feature representations for different fonts, text orientations, and complex scenes. Surprisingly, SSM shows superior performance for contextless text images, which may be attributed to the fact that embedding linguistic learning in visual space can avoid the text distribution dependence of explicit text correction.
@@ 4.4 Ablations and analysis
Effectiveness of architecture. The third row of Tab. 4 indicates that there is a 3.0% average performance improvement on Union14M benchmarks with the pixel-level reconstruction, which need the combined effect of prompt generator $\mathcal{G(\cdot)}$ transformer regressor $\mathcal{R(\cdot)}$ and pixel decoder $\mathcal{D(\cdot)}$. However, removing the regressor $\mathcal{R(\cdot)}$ results in a slight performance downgrade. The result demonstrates that the transformer regressor plays an indispensable role in SSM by guiding and disentangling the features of the original and inverted images. Based on the pixel reconstruction, the Joint use of the EMA mechanism and project module $\mathcal{H(\cdot)}$ further brings accuracy gains of 1.3% as it promotes learning discriminative character semantics and modeling spatial context at the semantic level. Here, $\mathcal{H(\cdot)}$ is mainly used to prevent excessive background noise in negative sample sampling and to map the decoupling feature to high-dimensional space. Lastly, the addition of Irregular view augmentations improves the average test accuracy up to 84.3% due to diverse perspectives.
Ablations on the types of superimposed input. We compared our superposition strategies with add noise &\&& blur and randomly add another text image. As shown in Tab. 5(a), the performance gains of add another text image are close to that of add noise &\&& blur, both of which are lower than existing superposition strategies. We attribute this to the fact that the model can easily reconstruct the two images with inconsistent text content and different background styles from low-level pixel differences, limiting the exploration of intrinsic character semantics and linguistic information. We also found that individually superposing a 180-degree rotation view is more effective than individually superposing a specific flip view. The result suggests that the rotation influences both character orientation and sequence reading order, helping capture learning richer character semantics and spatial context relationships during reconstruction. Finally, the best performance (84.3%) is achieved by combining HS, VS, and RS strategies, which improve the diversity of the superimposed input.
@@
Ablations on the feature-level reconstruction loss. The first and second rows of Tab. 5(b) shows that only using distance measure functions such as MSE $L_{mse}$ and Cosine $L_{cos}$ results in lower word accuracy (82.5% vs 94.3%). We attribute this to the fact that only minimizing the feature distance between predictions and targets may lead to feature collapse. Besides, Tab. 5(b) shows that our discriminative consistency loss $L_{dis}$ outperforms the dino loss $L_{dino}$ by 0.3%. The combined use of $L_{dis}$ and $L_{mse}$ ultimately yielded the best results. $L_{dis}$ can mitigate the feature collapse of $L_{mse}$ by reducing the similarity between negative samples. Besides, $L_{dis}$ focuses on enhancing discriminative features for character semantics while $L_{mse}$ is more sensitive to spatial structures, which helps model linguistic information in spatial context at the semantic level.